OpenStudy (anonymous):

What is the simplified form of y^2+y-6/y^2+2y-3?

OpenStudy (anonymous):

First you factor both of them. y^2 +y-6 = (y-2) (y+3) y^2+2y-3 = (y+3) (y-1) Then you cancel the same factors. This will leave you with (y-2)/(y-1) Hope it helped! :D

OpenStudy (anonymous):

hey i always write "i hope that helped :D" at the end of my explanation ... biter ^^

OpenStudy (anonymous):

Hahaha well, I just really hope it helped =))))) Fine, I'll think of my own tag line hahaha :P =)))

OpenStudy (anonymous):

im sorry for being mean .... i didn't mean it like that

OpenStudy (anonymous):

I'm kidding dude :))) And you weren't mean at all :))

OpenStudy (anonymous):

lol^^^^